OpenStudy (anonymous):

The length of the rectangle is twice the width. The area is 800 yd^2. Find the length and the width.

OpenStudy (anonymous):

width = x length = 2x area = 2x^2 2x^2 = 800 x^2 = 400 x = 20 width = 20 length = 40

OpenStudy (saifoo.khan):

x (2x) = 800 2x^2 = 800 x^2 = 400 x = 20

OpenStudy (saifoo.khan):

width = x = 20 length = 2x = 40

OpenStudy (saifoo.khan):

lol, the cheater must get ready for getting ban!